OpenStudy (chris215):

please help I will fan and medal! Find the coefficient of the squared term in the simplified form for the second derivative, f "(x) for f(x) = (x^3 + 3x^2 + 3)(3x^3 − 6x^2 − 8x + 1) . Use the hyphen symbol, -, for negative values.

OpenStudy (chris215):

i got 108?

OpenStudy (solomonzelman):

I don't think so.

OpenStudy (solomonzelman):

It might be a pain in the neck to expand it, but once you expand it to find the second derivative and then the coefficient of x^2 shouldn't be much trouble.

OpenStudy (solomonzelman):

Oh, you don't need to expand.

OpenStudy (michele_laino):

Hint: I got this expression for \(f(x)\): \(f(x)=3x^6+3x^5-26x^4-14x^3-15x^2-24x+3\)

OpenStudy (anonymous):

f"x=90x^4+60x^3-312x^2-84x-15

OpenStudy (anonymous):

That should be the answer

OpenStudy (anonymous):

after taking two derivatives of what Michele Lanio got

OpenStudy (solomonzelman):

You know that the second derivative of ax^4 will yield the coefficient you need. What is the ax^4 in this case? (x^3)•(-8x) + (-6x^2)•(3x^2) ______________ -8x^4-18x^4=-26x^4 g(x)=-26x^4 g''(x)=(-26•3•4)x^2

OpenStudy (solomonzelman):

You can ignore the rest of the terms in the f(x), just the products of those that give x^4.